# Ergodic behaviors in reversible 3-state cellular automata

### v1

@@ Appendix B Detailed properties of 3-state RCA
17348625 $\mathcal{C+T}^{E}$ IVa power 0.27 const 0.51(6) const 0.07(6) 0 0 0 0
46513287 $\mathcal{C+T}$ IVa power 0.275 const 0.51(6) const 0.07(6) 0 0 0 0
48617352 $\mathcal{C+T}$ IVa power 0.26(2) const 0.51(6) const 0.07(6) 0 0 0 0
13246578 $\mathcal{C+T}^{E}$ IVa power 0.286 const 0.51(6) const 0.07(6) 0 0 0 0
34681572 $\mathcal{CPT}^{*}$ IVb power 1.018 const 0.0 const 0.111 1 3 9 24 72
23657418 $\mathcal{CPT}^{*}$ IVb power 1.012 const 0.0 const 0.111 1 4 11 32 95
21654738 $\mathcal{CPT}^{*}$ IVb power 1.029 const 0.0 const 0.111 2 5 14 41 122
26458731 $\mathcal{CPT}^{*}$ IVb power 1.014 const 0.0 const 0.111 1 4 11 32 95
56428731 $\mathcal{CPT}^{*}$ IVb power 1.019 const 0.0 const 0.111 1 3 9 24 59
23651478 $\mathcal{CPT}$ IVb power 4.0(1) power 0.507 const 0.111 1 2 5 9 18
24657318 $\mathcal{CPT}$ IVb power 4.0(1) power 0.506 const 0.111 1 2 5 9 18
14375628 $\mathcal{CPT}$ IVb power 2.52(8) const 0.36(8) const 0.09(3) 1 6 26 136
45612738 $\mathcal{CPT}$ IVb power 2.78(7) const 0.17(8) const 0.222 3 18 111
45623187 $\mathcal{CPT}$ IVb power 2.57(6) const 0.36(8) const 0.09(3) 1 6 26 136
12645738 $\mathcal{CPT}$ IVb power 2.81(9) const 0.17(8) const 0.222 3 18 111
54628371 $\mathcal{CPT}^{*}$ IVb power 1.03(1) const 0.333 const 0.111 2 6 19 56 168
54623178 $\mathcal{CPT}^{*}$ IVb power 1.03(1) const 0.333 const 0.111 2 6 19 56 168
34781562 $\mathcal{CPT}^{*}$ IVb power 1.018 const 0.333 const 0.111 2 7 21 63 189
23857416 $\mathcal{CPT}^{*}$ IVb power 1.016 const 0.333 const 0.111 2 8 24 72 216
54621738 $\mathcal{CPT}^{*}$ IVb power 1.0(1) const 0.333 const 0.111 3 8 24 69 207
24658371 $\mathcal{CPT}^{*}$ IVb power 1.015 const 0.333 const 0.111 2 6 19 56 168
56428371 $\mathcal{CPT}^{*}$ IVb power 1.02(1) const 0.333 const 0.111 2 7 21 63 189
24753168 $\mathcal{CPT}^{*}$ IVb power 1.01(4) const 0.333 const 0.111 2 6 20 60 180
54728361 $\mathcal{CPT}^{*}$ IVb power 1.02(6) const 0.333 const 0.111 2 6 19 57 171
54723168 $\mathcal{CPT}^{*}$ IVb power 1.02(5) const 0.333 const 0.111 2 6 19 57 171
26458371 $\mathcal{CPT}^{*}$ IVb power 1.013 const 0.333 const 0.111 2 8 24 72 216
24758361 $\mathcal{CPT}^{*}$ IVb power 1.01(5) const 0.333 const 0.111 2 6 20 60 180
24651873 $\mathcal{CPT}^{*}$ IVb power 1.0(1) const 0.333 const 0.111 3 8 24 69 207
24653178 $\mathcal{CPT}^{*}$ IVb power 1.01 const 0.333 const 0.111 2 6 19 56 168
24651738 $\mathcal{CPT}^{*}$ IVb power 1.024 const 0.0 const 0.111 2 5 14 41 122
14325678 $\mathcal{CP+T}$ IVb power 1.24(3) const 0.48(6) const 0.12(3) 3 19 149
45621387 $\mathcal{CT+P}$ IVb power 1.23(3) const 0.48(6) const 0.12(3) 3 19 149 1181
12345687 $\mathcal{C+P+T}$ IVb power 2.56(8) const 0.23(8) const 0.222 4 21 145
12645378 $\mathcal{C+P+T}$ IVb power 2.6(9) const 0.23(8) const 0.222 4 21 145
45312687 $\mathcal{C+P+T}$ IVb power 2.57(8) const 0.23(8) const 0.222 4 21 145
45612378 $\mathcal{C+P+T}$ IVb power 2.6(8) const 0.23(8) const 0.222 4 21 145
21654387 $\mathcal{C+P+T}$ IVb power 1.995 power 0.502 const 0.111 3 5 9 17 33
54321678 $\mathcal{C+P+T}$ IVb power 1.99(1) power 0.501 const 0.111 3 5 9 17 33
54621387 $\mathcal{C+P+T}$ IVb power 1.99(1) power 0.501 const 0.111 3 5 9 17 33
21354678 $\mathcal{C+P+T}$ IVb power 2.0(1) power 0.502 const 0.111 3 5 9 17 33
21354687 $\mathcal{C+P+T}^{*}$ IVb power 1.033 const 0.333 const 0.111 4 12 36 108 324
54621378 $\mathcal{C+P+T}^{*}$ IVb power 1.033 const 0.333 const 0.111 4 12 36 108 324
54321687 $\mathcal{C+P+T}^{*}$ IVb power 1.033 const 0.0 const 0.111 2 6 18 54 162
21654378 $\mathcal{C+P+T}^{*}$ IVb power 1.033 const 0.0 const 0.111 2 6 18 54 162
@@
15342687 $\mathcal{C+T}$ IVb power 2.54(8) const 0.1(9) const 0.222 3 18 111
@@
15642378 $\mathcal{C+T}$ IVb power 2.49(7) const 0.1(9) const 0.222 3 18 111
@@
21754836 $\mathcal{C+T}$ IVb power 2.01(1) const 0.0 const 0.111 2 4 8 16 32
@@
54721836 $\mathcal{C+T}$ IVb power 2.007 exp 0.307 const 0.111 2 4 8 16 32
@@
26534187 $\mathcal{CP+PT}$ IVb power 2.54(6) const 0.5(6) const 0.07(5) 4 21 145
25614387 $\mathcal{CP+PT}$ IVb power 2.57(7) const 0.56(2) const 0.11(2) 4 21 145
23164578 $\mathcal{CP+PT}$ IVb power 3.1(1) power 0.496 power 0.47(1) 1 3 4 9 17
25314687 $\mathcal{CP+PT}$ IVb power 1.995 const 0.25 const 0.028 3 5 9 17 33
@@
35216487 $\mathcal{CP+PT}$ IVb power 1.997 const 0.25 const 0.028 3 5 9 17 33
23164587 $\mathcal{CP+PT}^{*}$ IVb power 1.033 const 0.25 const 0.028 2 6 18 54 162
21378645 $\mathcal{P+T}$ IVb power 2.59(7) const 0.5(6) const 0.07(5) 4 21 145 1077
21387654 $\mathcal{P+T}^{*}$ IVb power 1.033 const 0.25 const 0.028 2 6 18 54 162
21687354 $\mathcal{P+T}$ IVb power 3.1(1) power 0.496 power 0.46(1) 1 3 5 9 17
12354678 $\mathcal{P+T}$ IVb power 2.57(8) const 0.56(2) const 0.11(2) 4 21 145 1077
12354687 $\mathcal{P+T}$ IVb power 2.0(1) const 0.25 const 0.028 3 5 9 17 33
@@
12387654 $\mathcal{P+T}$ IVb power 2.004 const 0.25 const 0.028 3 5 9 17 33
12745863 $\mathcal{PT+C}$ IVb power 2.54(5) const 0.13(8) const 0.222 3 18 111
21754863 $\mathcal{PT+C}^{*}$ IVb power 1.033 const 0.333 const 0.111 3 9 27 81 243
23156478 $\mathcal{PT+C}$ IVb power 3.0(2) power 0.496 const 0.111 1 4 7 13 25
23156487 $\mathcal{PT+C}^{*}$ IVb power 1.04(1) const 0.333 const 0.111 2 7 21 63 189
24351678 $\mathcal{PT+C}$ IVb power 2.0(3) const 0.0 const 0.111 2 4 8 16 32
24351687 $\mathcal{PT+C}^{*}$ IVb power 1.033 const 0.333 const 0.111 3 9 27 81 243
24651378 $\mathcal{PT+C}^{*}$ IVb power 1.033 const 0.333 const 0.111 3 9 27 81 243
24651387 $\mathcal{PT+C}$ IVb power 2.0(3) const 0.0 const 0.111 2 4 8 16 32
24751863 $\mathcal{PT+C}^{*}$ IVb power 1.033 const 0.333 const 0.111 4 12 36 108 324
24851736 $\mathcal{PT+C}^{*}$ IVb power 1.033 const 0.0 const 0.111 2 6 18 54 162
26453178 $\mathcal{PT+C}^{*}$ IVb power 1.0(2) const 0.333 const 0.111 2 8 24 72 216
26453187 $\mathcal{PT+C}$ IVb power 3.0(2) power 0.499 const 0.111 2 4 8 16
27358641 $\mathcal{PT+C}$ IVb power 3.0(2) power 0.496 const 0.111 2 4 8 16 32
27658341 $\mathcal{PT+C}^{*}$ IVb power 1.0(2) const 0.333 const 0.111 2 8 24 72 216
28357614 $\mathcal{PT+C}^{*}$ IVb power 1.038 const 0.333 const 0.111 2 7 21 63 189
28657314 $\mathcal{PT+C}$ IVb power 3.0(2) power 0.495 const 0.111 1 4 7 13 25
34261578 $\mathcal{PT+C}$ IVb power 3.0(2) power 0.498 const 0.111 2 4 8 16 32
34261587 $\mathcal{PT+C}^{*}$ IVb power 1.0(2) const 0.333 const 0.111 2 8 24 72 216
45712863 $\mathcal{PT+C}$ IVb power 2.56(6) const 0.13(8) const 0.222 3 18 111
54721863 $\mathcal{PT+C}^{*}$ IVb power 1.033 const 0.333 const 0.111 3 9 27 81 243
56423178 $\mathcal{PT+C}^{*}$ IVb power 1.03(1) const 0.333 const 0.111 2 7 21 63 189
56423187 $\mathcal{PT+C}$ IVb power 3.0(2) power 0.497 const 0.111 1 4 7 13 25
57328641 $\mathcal{PT+C}$ IVb power 3.0(2) power 0.497 const 0.111 1 4 5 13 25
57628341 $\mathcal{PT+C}^{*}$ IVb power 1.04(1) const 0.333 const 0.111 2 7 21 63 189
58327614 $\mathcal{PT+C}^{*}$ IVb power 1.0(2) const 0.333 const 0.111 2 8 24 72 216
58627314 $\mathcal{PT+C}$ IVb power 3.0(2) power 0.498 const 0.111 2 4 8 16 32
38671524 $\mathcal{CPT}$ exp 0.9(3) power 0.514 ? ? 0 0 1 1 1
36427518 $\mathcal{CPT}$ exp 0.9(3) power 0.514 ? ? 0 0 1 1 1
23651487 $\mathcal{CPT}$ exp 0.45(2) ? ? const 0.111 1 2 4 8 16
24357618 $\mathcal{CPT}$ exp 0.45(2) ? ? const 0.111 1 2 4 8 16
34728561 $\mathcal{CPT}$ exp 0.54(1) exp 0.048 ? ? 0 0 0 0 0
14785362 $\mathcal{CPT}$ exp 0.55(1) exp 0.138 ? ? 0 0 0 0 0
45327618 $\mathcal{CPT}$ exp 0.55(1) exp 0.159 ? ? 0 0 0 0 0
36412578 $\mathcal{CPT}$ exp 0.54(1) exp 0.14 ? ? 0 0 0 0 0
34712568 $\mathcal{CPT}$ exp 0.54(1) exp 0.138 ? ? 0 0 0 0 0
13625478 $\mathcal{CPT}$ exp 0.53(1) exp 0.054 ? ? 0 0 0 0 0
34627518 $\mathcal{CPT}$ exp 0.54(2) power 0.53 ? ? 0 1 1 1 1
24653187 $\mathcal{CPT}$ exp 0.59(4) ? ? const 0.111 1 2 4 8 16
24358671 $\mathcal{CPT}$ exp 0.6(5) ? ? const 0.111 1 2 4 8 16
23564178 $\mathcal{CP}$ exp 0.98(9) ? ? ? ? 0 0 0 0 0 0
18647352 $\mathcal{C+T}$ 0 0 0 0 0 0
16543287 $\mathcal{C+T}$ 0 0 0 0 0 0
35162478 $\mathcal{C+T}$ 0 0 0 0 0 0
17648325 $\mathcal{C+T}$ exp 0.546 exp 0.406 ? ? 0 0 0 0 0 0
46513278 $\mathcal{C+T}$ exp 0.55 exp 0.404 ? ? 0 0 0 0 0 0
47318625 $\mathcal{C+T}^{E}$ const 0.0 0 0 0 0 0 0
48317652 $\mathcal{C+T}$ exp 0.549 exp 0.402 ? ? 0 0 0 0 0 0
68437152 $\mathcal{C+T}$ exp 0.55(4) ? ? ? ? 0 0 0 0 0
37168425 $\mathcal{C+T}$ exp 0.55(4) ? ? ? ? 0 0 0 0 0 0
13246587 $\mathcal{C+T}$ exp 0.552 exp 0.405 ? ? 0 0 0 0 0 0
12678345 $\mathcal{P+T}$ exp 0.81(9) power 0.405 ? ? 1 1 1 1 1
12378654 $\mathcal{P}$ exp 0.89(9) ? ? ? ? 1 1 1 1 1
21678354 $\mathcal{P}$ exp 0.98(9) ? ? ? ? 0 0 0 0 0
45378612 $\mathcal{P}$ exp 0.56(1) power 0.526 ? ? 0 0 0 0 0
38267514 $\mathcal{PT+C}^{*}$ exp 0.172 exp 0.403 ? ? 0 2 4 10 22 54
67538241 $\mathcal{PT+C}^{*}$ exp 0.165 exp 0.398 ? ? 0 2 4 10 22 54
Appendix C Unclassifiable rules
@@
The first examples of these rules have their return times increase exponentially to a plateau. Known representatives are in the East class: rules 12748536, 17245836, 13745826, 12746835. It is unclear what mechanism can cause the orbit length to plateau. Secondly, we find cases where the orbit length sporadically differs by orders of magnitude for small changes of the system size $L$, e.g., in models with rules 43216578, 13746825, 17248536, 47318625. They also exhibit a correlation function which decays in 1 step. Some of these could be explained as perfect maps [17]. One example from each of the two odd groups is presented in Figure 34.
Figure 34: a) Dynamics of the model with rule 12748536, b) Dynamics of the model with rule 43216578, c) Orbit lengths as a function of the system size.

### v2

@@ Appendix B Detailed properties of 3-state RCA
17348625 $\mathcal{C+T}^{E}$ IVa poly 0.27 const 0.51(6) const 0.07(6) 0 0 0 0
46513287 $\mathcal{C+T}$ IVa poly 0.275 const 0.51(6) const 0.07(6) 0 0 0 0
48617352 $\mathcal{C+T}$ IVa poly 0.26(2) const 0.51(6) const 0.07(6) 0 0 0 0
13246578 $\mathcal{C+T}^{E}$ IVa poly 0.286 const 0.51(6) const 0.07(6) 0 0 0 0
34681572 $\mathcal{CPT}^{*}$ IVb poly 1.018 const 0.0 const 0.111 1 3 9 24 72
23657418 $\mathcal{CPT}^{*}$ IVb poly 1.012 const 0.0 const 0.111 1 4 11 32 95
21654738 $\mathcal{CPT}^{*}$ IVb poly 1.029 const 0.0 const 0.111 2 5 14 41 122
26458731 $\mathcal{CPT}^{*}$ IVb poly 1.014 const 0.0 const 0.111 1 4 11 32 95
56428731 $\mathcal{CPT}^{*}$ IVb poly 1.019 const 0.0 const 0.111 1 3 9 24 59
23651478 $\mathcal{CPT}$ IVb poly 4.0(1) poly 0.507 const 0.111 1 2 5 9 18
24657318 $\mathcal{CPT}$ IVb poly 4.0(1) poly 0.506 const 0.111 1 2 5 9 18
14375628 $\mathcal{CPT}$ IVb poly 2.52(8) const 0.36(8) const 0.09(3) 1 6 26 136
45612738 $\mathcal{CPT}$ IVb poly 2.78(7) const 0.17(8) const 0.222 3 18 111
45623187 $\mathcal{CPT}$ IVb poly 2.57(6) const 0.36(8) const 0.09(3) 1 6 26 136
12645738 $\mathcal{CPT}$ IVb poly 2.81(9) const 0.17(8) const 0.222 3 18 111
54628371 $\mathcal{CPT}^{*}$ IVb poly 1.03(1) const 0.333 const 0.111 2 6 19 56 168
54623178 $\mathcal{CPT}^{*}$ IVb poly 1.03(1) const 0.333 const 0.111 2 6 19 56 168
34781562 $\mathcal{CPT}^{*}$ IVb poly 1.018 const 0.333 const 0.111 2 7 21 63 189
23857416 $\mathcal{CPT}^{*}$ IVb poly 1.016 const 0.333 const 0.111 2 8 24 72 216
54621738 $\mathcal{CPT}^{*}$ IVb poly 1.0(1) const 0.333 const 0.111 3 8 24 69 207
24658371 $\mathcal{CPT}^{*}$ IVb poly 1.015 const 0.333 const 0.111 2 6 19 56 168
56428371 $\mathcal{CPT}^{*}$ IVb poly 1.02(1) const 0.333 const 0.111 2 7 21 63 189
24753168 $\mathcal{CPT}^{*}$ IVb poly 1.01(4) const 0.333 const 0.111 2 6 20 60 180
54728361 $\mathcal{CPT}^{*}$ IVb poly 1.02(6) const 0.333 const 0.111 2 6 19 57 171
54723168 $\mathcal{CPT}^{*}$ IVb poly 1.02(5) const 0.333 const 0.111 2 6 19 57 171
26458371 $\mathcal{CPT}^{*}$ IVb poly 1.013 const 0.333 const 0.111 2 8 24 72 216
24758361 $\mathcal{CPT}^{*}$ IVb poly 1.01(5) const 0.333 const 0.111 2 6 20 60 180
24651873 $\mathcal{CPT}^{*}$ IVb poly 1.0(1) const 0.333 const 0.111 3 8 24 69 207
24653178 $\mathcal{CPT}^{*}$ IVb poly 1.01 const 0.333 const 0.111 2 6 19 56 168
24651738 $\mathcal{CPT}^{*}$ IVb poly 1.024 const 0.0 const 0.111 2 5 14 41 122
14325678 $\mathcal{CP+T}$ IVb poly 1.24(3) const 0.48(6) const 0.12(3) 3 19 149
45621387 $\mathcal{CT+P}$ IVb poly 1.23(3) const 0.48(6) const 0.12(3) 3 19 149 1181
12345687 $\mathcal{C+P+T}$ IVb poly 2.56(8) const 0.23(8) const 0.222 4 21 145
12645378 $\mathcal{C+P+T}$ IVb poly 2.6(9) const 0.23(8) const 0.222 4 21 145
45312687 $\mathcal{C+P+T}$ IVb poly 2.57(8) const 0.23(8) const 0.222 4 21 145
45612378 $\mathcal{C+P+T}$ IVb poly 2.6(8) const 0.23(8) const 0.222 4 21 145
21654387 $\mathcal{C+P+T}$ IVb poly 1.995 poly 0.502 const 0.111 3 5 9 17 33
54321678 $\mathcal{C+P+T}$ IVb poly 1.99(1) poly 0.501 const 0.111 3 5 9 17 33
54621387 $\mathcal{C+P+T}$ IVb poly 1.99(1) poly 0.501 const 0.111 3 5 9 17 33
21354678 $\mathcal{C+P+T}$ IVb poly 2.0(1) poly 0.502 const 0.111 3 5 9 17 33
21354687 $\mathcal{C+P+T}^{*}$ IVb poly 1.033 const 0.333 const 0.111 4 12 36 108 324
54621378 $\mathcal{C+P+T}^{*}$ IVb poly 1.033 const 0.333 const 0.111 4 12 36 108 324
54321687 $\mathcal{C+P+T}^{*}$ IVb poly 1.033 const 0.0 const 0.111 2 6 18 54 162
21654378 $\mathcal{C+P+T}^{*}$ IVb poly 1.033 const 0.0 const 0.111 2 6 18 54 162
@@
15342687 $\mathcal{C+T}$ IVb poly 2.54(8) const 0.1(9) const 0.222 3 18 111
@@
15642378 $\mathcal{C+T}$ IVb poly 2.49(7) const 0.1(9) const 0.222 3 18 111
@@
21754836 $\mathcal{C+T}$ IVb poly 2.01(1) const 0.0 const 0.111 2 4 8 16 32
@@
54721836 $\mathcal{C+T}$ IVb poly 2.007 exp 0.307 const 0.111 2 4 8 16 32
@@
26534187 $\mathcal{CP+PT}$ IVb poly 2.54(6) const 0.5(6) const 0.07(5) 4 21 145
25614387 $\mathcal{CP+PT}$ IVb poly 2.57(7) const 0.56(2) const 0.11(2) 4 21 145
23164578 $\mathcal{CP+PT}$ IVb poly 3.1(1) poly 0.496 poly 0.47(1) 1 3 4 9 17
25314687 $\mathcal{CP+PT}$ IVb poly 1.995 const 0.25 const 0.028 3 5 9 17 33
@@
35216487 $\mathcal{CP+PT}$ IVb poly 1.997 const 0.25 const 0.028 3 5 9 17 33
23164587 $\mathcal{CP+PT}^{*}$ IVb poly 1.033 const 0.25 const 0.028 2 6 18 54 162
21378645 $\mathcal{P+T}$ IVb poly 2.59(7) const 0.5(6) const 0.07(5) 4 21 145 1077
21387654 $\mathcal{P+T}^{*}$ IVb poly 1.033 const 0.25 const 0.028 2 6 18 54 162
21687354 $\mathcal{P+T}$ IVb poly 3.1(1) poly 0.496 poly 0.46(1) 1 3 5 9 17
12354678 $\mathcal{P+T}$ IVb poly 2.57(8) const 0.56(2) const 0.11(2) 4 21 145 1077
12354687 $\mathcal{P+T}$ IVb poly 2.0(1) const 0.25 const 0.028 3 5 9 17 33
@@
12387654 $\mathcal{P+T}$ IVb poly 2.004 const 0.25 const 0.028 3 5 9 17 33
12745863 $\mathcal{PT+C}$ IVb poly 2.54(5) const 0.13(8) const 0.222 3 18 111
21754863 $\mathcal{PT+C}^{*}$ IVb poly 1.033 const 0.333 const 0.111 3 9 27 81 243
23156478 $\mathcal{PT+C}$ IVb poly 3.0(2) poly 0.496 const 0.111 1 4 7 13 25
23156487 $\mathcal{PT+C}^{*}$ IVb poly 1.04(1) const 0.333 const 0.111 2 7 21 63 189
24351678 $\mathcal{PT+C}$ IVb poly 2.0(3) const 0.0 const 0.111 2 4 8 16 32
24351687 $\mathcal{PT+C}^{*}$ IVb poly 1.033 const 0.333 const 0.111 3 9 27 81 243
24651378 $\mathcal{PT+C}^{*}$ IVb poly 1.033 const 0.333 const 0.111 3 9 27 81 243
24651387 $\mathcal{PT+C}$ IVb poly 2.0(3) const 0.0 const 0.111 2 4 8 16 32
24751863 $\mathcal{PT+C}^{*}$ IVb poly 1.033 const 0.333 const 0.111 4 12 36 108 324
24851736 $\mathcal{PT+C}^{*}$ IVb poly 1.033 const 0.0 const 0.111 2 6 18 54 162
26453178 $\mathcal{PT+C}^{*}$ IVb poly 1.0(2) const 0.333 const 0.111 2 8 24 72 216
26453187 $\mathcal{PT+C}$ IVb poly 3.0(2) poly 0.499 const 0.111 2 4 8 16
27358641 $\mathcal{PT+C}$ IVb poly 3.0(2) poly 0.496 const 0.111 2 4 8 16 32
27658341 $\mathcal{PT+C}^{*}$ IVb poly 1.0(2) const 0.333 const 0.111 2 8 24 72 216
28357614 $\mathcal{PT+C}^{*}$ IVb poly 1.038 const 0.333 const 0.111 2 7 21 63 189
28657314 $\mathcal{PT+C}$ IVb poly 3.0(2) poly 0.495 const 0.111 1 4 7 13 25
34261578 $\mathcal{PT+C}$ IVb poly 3.0(2) poly 0.498 const 0.111 2 4 8 16 32
34261587 $\mathcal{PT+C}^{*}$ IVb poly 1.0(2) const 0.333 const 0.111 2 8 24 72 216
45712863 $\mathcal{PT+C}$ IVb poly 2.56(6) const 0.13(8) const 0.222 3 18 111
54721863 $\mathcal{PT+C}^{*}$ IVb poly 1.033 const 0.333 const 0.111 3 9 27 81 243
56423178 $\mathcal{PT+C}^{*}$ IVb poly 1.03(1) const 0.333 const 0.111 2 7 21 63 189
56423187 $\mathcal{PT+C}$ IVb poly 3.0(2) poly 0.497 const 0.111 1 4 7 13 25
57328641 $\mathcal{PT+C}$ IVb poly 3.0(2) poly 0.497 const 0.111 1 4 5 13 25
57628341 $\mathcal{PT+C}^{*}$ IVb poly 1.04(1) const 0.333 const 0.111 2 7 21 63 189
58327614 $\mathcal{PT+C}^{*}$ IVb poly 1.0(2) const 0.333 const 0.111 2 8 24 72 216
58627314 $\mathcal{PT+C}$ IVb poly 3.0(2) poly 0.498 const 0.111 2 4 8 16 32
38671524 $\mathcal{CPT}$ ? exp 0.9(3) poly 0.514 ? ? 0 0 1 1 1
36427518 $\mathcal{CPT}$ ? exp 0.9(3) poly 0.514 ? ? 0 0 1 1 1
23651487 $\mathcal{CPT}$ ? exp 0.45(2) ? ? const 0.111 1 2 4 8 16
24357618 $\mathcal{CPT}$ ? exp 0.45(2) ? ? const 0.111 1 2 4 8 16
34728561 $\mathcal{CPT}$ ? exp 0.54(1) exp 0.048 ? ? 0 0 0 0 0
14785362 $\mathcal{CPT}$ ? exp 0.55(1) exp 0.138 ? ? 0 0 0 0 0
45327618 $\mathcal{CPT}$ ? exp 0.55(1) exp 0.159 ? ? 0 0 0 0 0
36412578 $\mathcal{CPT}$ ? exp 0.54(1) exp 0.14 ? ? 0 0 0 0 0
34712568 $\mathcal{CPT}$ ? exp 0.54(1) exp 0.138 ? ? 0 0 0 0 0
13625478 $\mathcal{CPT}$ ? exp 0.53(1) exp 0.054 ? ? 0 0 0 0 0
34627518 $\mathcal{CPT}$ ? exp 0.54(2) poly 0.53 ? ? 0 1 1 1 1
24653187 $\mathcal{CPT}$ ? exp 0.59(4) ? ? const 0.111 1 2 4 8 16
24358671 $\mathcal{CPT}$ ? exp 0.6(5) ? ? const 0.111 1 2 4 8 16
23564178 $\mathcal{CP}$ ? exp 0.98(9) ? ? ? ? 0 0 0 0 0 0
18647352 $\mathcal{C+T}$ ? ? ? ? ? ? ? 0 0 0 0 0 0
16543287 $\mathcal{C+T}$ ? ? ? ? ? ? ? 0 0 0 0 0 0
35162478 $\mathcal{C+T}$ ? ? ? ? ? ? ? 0 0 0 0 0 0
17648325 $\mathcal{C+T}$ ? exp 0.546 exp 0.406 ? ? 0 0 0 0 0 0
46513278 $\mathcal{C+T}$ ? exp 0.55 exp 0.404 ? ? 0 0 0 0 0 0
47318625 $\mathcal{C+T}^{E}$ ? ? ? ? ? const 0.0 0 0 0 0 0 0
48317652 $\mathcal{C+T}$ ? exp 0.549 exp 0.402 ? ? 0 0 0 0 0 0
68437152 $\mathcal{C+T}$ ? exp 0.55(4) ? ? ? ? 0 0 0 0 0
37168425 $\mathcal{C+T}$ ? exp 0.55(4) ? ? ? ? 0 0 0 0 0 0
13246587 $\mathcal{C+T}$ ? exp 0.552 exp 0.405 ? ? 0 0 0 0 0 0
12678345 $\mathcal{P+T}$ ? exp 0.81(9) poly 0.405 ? ? 1 1 1 1 1
12378654 $\mathcal{P}$ ? exp 0.89(9) ? ? ? ? 1 1 1 1 1
21678354 $\mathcal{P}$ ? exp 0.98(9) ? ? ? ? 0 0 0 0 0
45378612 $\mathcal{P}$ ? exp 0.56(1) poly 0.526 ? ? 0 0 0 0 0
38267514 $\mathcal{PT+C}^{*}$ ? exp 0.172 exp 0.403 ? ? 0 2 4 10 22 54
67538241 $\mathcal{PT+C}^{*}$ ? exp 0.165 exp 0.398 ? ? 0 2 4 10 22 54
@@
The first examples of these rules have their return times increase exponentially to a plateau. It is unclear what mechanism can cause the orbit length to plateau. Secondly, we find cases where the orbit length sporadically differs by orders of magnitude for small changes of the system size $L$, e.g., in models with rules 43216578, 13746825, 17248536, 47318625. They also exhibit a correlation function which decays in 1 step. Some of these could be explained as perfect maps [17].
@@
Appendix C Extended simulations of subdiffusive models
A select few subdiffusive models exhibit slowly converging dynamical exponents $z$. To rule out the emergence of any new exotic universality classes, and to confirm the dynamical exponents on asymptotic scales, we performed additional simulations for systems of size $2^{20}$ and times up to $2^{21}$. This was done for the following set of rules:
  1. 1.
    46532178,
  2. 2.
    16435278 (and its equivalent 45678321),
  3. 3.
    14625387 (and its equivalent 45321678),
  4. 4.
    25614378,
  5. 5.
    12654378 (equivalent to 25614378).
We computed the correlators and examined the convergence of the dynamical exponents as a function of time. The results are presented in Figure 36. It appears that almost all dynamical exponents converge rather slowly to $z=4$. The exception are the cellular automata with rules 14625387 and 45321678, for which the dynamical exponents of the $\langle[\varnothing]_{xt}[\varnothing]_{00}\rangle$ correlators did not converge even at times $2^{21}$. Nevertheless, their values appear to be approaching $z=4$ as well. When available, this long-time data is used in the classification table for $z$ instead of the data extracted from the shorter-time correlators computed for other models.
Figure 36: Correlators $\hat{C}(x,t)$ and $\hat{D}(x,t)$ computed for longer timescales for subdiffusive models. The blue and orange points correspond to $\hat{C}(x,t)$ and $\hat{D}(x,t)$. respectively, while the dashed lines denote power-law fits of the form $t^{-1/z}$.